# Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees

### v2

@@ IV Simulation Results
We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}$. For comparison, relaxed and robust models from [RobustNawaf] are used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The chosen element battery parameters are obtained from a Tesla Powerwall and we set $N=1000$ and $M=1$. The results from using each model are shown in Fig. 3.
Figure 3: Reference tracking optimization.
The RCB model and relaxed battery model produced similar optimal dispatch. However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate, resulting in the sudden drop in power seen in Fig. 3. The MILP, robust, and RCB models produced realizable results, with the robust formulation resulting in worse tracking performance. The MILP and RCB formulations produced comparable results, however, with a significant computational advantage for the RCB model.
@@ V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems.

### v4

@@ IV Simulation Results
We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The battery element parameters given in Table I were used. The two use-cases are a) power reference tracking and b) revenue maximization.
IV-A Power Reference Tracking
In this use case, the goal is to minimize the power tracking error for a given reference signal as shown below,
$\min_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}.$ (14)
A timestep of $\Delta t=5$ min, along with $N=100$ was used. The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively. The results from using each model are shown in Fig. 5 and the solve times along with the mean squared tracking error are shown in Table II.
Figure 5: Open-loop implementation of power optimal reference tracking power schedules using MILP, relaxed, robust and proposed RCB models.
TABLE II: Power Reference Tracking Performance Metrics
Model Solver Predicted Actual
Time (ms) MSE (kW2) MSE (kW2)
MILP 50,140 0 0
Relaxed 30 0 1841
Robust 34 2.12 2.12
RCB (M = 1) 30 0.72 0.72
RCB (M = 5) 30 0.013 0.013
RCB (M = 10) 30 0 0
The MILP model yielded the optimal solution and was able to track the reference signal with no tracking error. However, the MILP also came with a significant computational burden compared to the rest of the formulations. The other formulations are linear and, as a result, have significantly faster solve times. The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs leading to a large realized tracking error. The robust model yields a realizable yet sub-optimal solution. The RCB model shows improved tracking error with increasing values of $M$ but even with $M=1$ the tracking error surpasses the robust model.
IV-B Revenue Maximization
In this use case, the goal is to maximize the predicted revenue from the composite battery system given a forecasted price signal. The objective function is given by,
$\max_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K}C[k](P_{\text{d}}[k]-P_{\text{c}}[k]),$ (15)
where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices for 24 hours, plotted in Fig 6.
Figure 6: CAISO day-ahead price forecasts
A timestep of $\Delta t=15$ min, along with $N=100$ was used. The RCB model was tested for $M=$ 1, 5, and 10 which correspond to $\delta t=$ 15 min, 3 min and 1.5 min respectively. An additional case was added with $M=900$ ($\delta t=1s$) to gain insight on model performance as $M\rightarrow\infty.$The results for the different models are shown in Table III.
TABLE III: Revenue Maximization Performance metrics
Model Solver Predicted Actual
Time (ms) Revenue (\$) Revenue (\$)
MILP 1250 2088.97 2088.97
Relaxed 1 2092.80 1940.35
Robust 1 1866.42 1866.42
RCB (M = 1) 1 1411.53 1411.53
RCB (M = 5) 1 1955.76 1955.76
RCB (M = 10) 1 2022.87 2022.87
RCB (M = 900) 1 2089.23 2089.23
Similar to the previous use-case, the MILP model produced the highest revenue but at a significantly larger computational cost. The relaxed model predicts the highest revenue but underestimates the SOC, leading to unpredicted battery saturation and a realized performance that is significantly worse. The robust model yields a realizable yet suboptimal solution. The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set. However, for f $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and approaches the predicted performance for larger values of $M$. It was also observed that the RCB surpasses the robust model for $M\geq 3$ ($\delta t\leq 5$ min).
@@ V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems or aggregate storage representations in energy system optimization models.